Circular No. 7332

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999gk IN NGC 4653 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report the serendipitous discovery by P. Berlind of a supernova in NGC 4653, during the course of spectroscopic observations with the F. L. Whipple Observatory (FLWO) 1.5-m telescope (+ FAST spectrograph). A spectrum of SN 1999gk, taken by Berlind on Dec. 15.53 UT, shows it to be a type-II supernova just after maximum light. The spectrum exhibits a blue continuum with strong P-Cyg Balmer lines and Fe II absorption. The Hα profile is quite peculiar, with the P-Cyg absorption showing two minima at expansion velocites of 16 100 and 8300 km/s, perhaps implying an asymmetric expansion and the possibility of strong polarization. Superimposed narrow Hα emission yields a recession velocity of 2490 km/s for the host galaxy. Concurrent images, taken by C. Hergenrother with the FLWO 1.2-m telescope (+ 4Shooter), yield a precise position for the supernova of α = 12^{h}43^{m}51^{s}.91, δ = −0^{o}32′55′′.4 (equinox 2000.0), which is 45′′.6 north and 15′′.3 east of the host galaxy, and magnitudes for the supernova V = 15.7 and R = 15.8. Coincidentally, the supernova is in the Landolt standard star field SA104.

GRB 991208 A. J. Castro-Tirado, Laboratorio de Astrofísica Espacial y Física Fun-

damental del Instituto Nacional de Técnica Aerospacial, Madrid, and Instituto de Astrofísica de Andalucía, Granada, on behalf of a large European collaboration; C. Blake and J. Wall, University of Oxford; and G. Feulner and U. Hopp, Universitäts Sternwarte, Hamburg, report: “We have imaged the 14 arcmin^{2} error box derived by the Ulysses, KONUS, and NEAR spacecraft for the extremely intense γ-ray burst detected on Dec. 8.19227 UT (Hurley et al., GCN 450). The images were taken with the 2.5-m Isaac Newton telescope (INT) at La Palma on Dec. 10.27 (I band) and at the 2.5m INT and 2.2-m telescope of the German-Spanish Calar Alto Observatory on Dec. 11.27 (R- and I-band filters). After a visual comparison with the Digital Sky Survey, a new source is clearly detected with R = 19.5 ± 0.1 (Dec. 11.27). Its position is α = 16^{h}33^{m}53^{s}.51, δ = +46^{o}27′21′′.5 (equinox 2000.0; ± 1′′). This object coincides with the previously unknown radio source reported by Frail et al. at position end figures 53^{s}.50, 20′′.9 (GCN 451) and therefore is taken to be the optical afterglow to GRB 991208. Further multiwavelength imaging and spectroscopy are encouraged. A finding

chart is posted at http://www.laeff.esa.es/~ajct/GRBs/GRB991208.”

1999 December 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green